OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2

OpenStudy (anonymous):

Let y=-2x+1 Then substitute the value of y to 2nd equation 4x+2(-2x+1)=1 4x-2x+1=1 0=0 It has infinite solution

OpenStudy (texaschic101):

2x + y = 1 ---> y = 1 - 2x now sub 1 - 2x in for y in the second equation... 4x + 2y = -2 4x + 2(1 - 2x) = - 2 now distribute the 2 through the parenthesis. Combine like terms. Get the x's on one side and everything else on the other side. Solve for x. Then sub what you got for x into either of the original equations to solve for y. Once you know x and y, you can check it by subbing in your known variables into either of the equations, and if they come out equal, then your answer is correct. Can you take it from here ?

OpenStudy (anonymous):

Ooops it's 4x+2(-2x+1)=-2 4x-4x+2=-2 0=0